### v1

# Determination of nuclear quadrupole moments of 25Mg, 87Sr, and 135,137Ba via configuration-interaction plus coupled-cluster approach

Abstract
[…]
I Introduction
The nuclear electric quadrupole moment $Q$ is an important parameter offering fundamental information on nuclear structure symmetry. Accurate knowledge of this parameter is important in multiple scientific fields like atomic, nuclear, and condensed matter physics, as well as chemistry and biology Qi et al. (2023, 2025); Guan et al. (2025); Heyde and Wood (2011); Campbell et al. (2016); Sinitsyn et al. (2012); Pyykkö (2018); Wong et al. (2006). An effective way to determine nuclear quadrupole moments is to integrate the experimentally-measured electric quadrupole hyperfine-structure constants ($B=234.9648867qQ$, in MHz), with the theoretically-calculated electric-field gradients $q$ (in $a.u.$). This method is independent of nuclear models and is one of the most accurate methods for extracting the nuclear quadrupole moments of unstable nuclei and heavy nuclei Sahoo et al. (2013); Bieroń et al. (2018); Li et al. (2021); Porsev et al. (2021); Papoulia et al. (2021); Skripnikov et al. (2021); Skripnikov and Barzakh (2024); Lu and Chang (2024); Zhang et al. (2025); Tang (2025). Its accuracy in extracting the nuclear quadrupole moment depends on calculated electric-field gradient as hyperfine spectra measured by spectral technology are high nowadays. Precise electric-field gradient calculation requires considering relativistic and electron correlation effects. Since the relativistic effect can be easily dealt with by solving the Dirac-Fock equation, electron correlation is the decisive factor for accurate electric-field gradients.
Recently, we developed a relativistic hybrid-method-based code for accurately calculating divalent atomic system properties. The hybrid method integrates the advantages of the configuration interaction method and the coupled-cluster method, and can simultaneously account for core-core, core-valence, and valence-valence correlations. We used it to calculate electric-field gradients of low-lying states in $^{43}\mathrm{Ca}$ and got an accurate nuclear quadrupole moment of $^{43}\mathrm{Ca}$ Tang (2025). In the present work, we apply the same method to obtain nuclear quadrupole moments of $^{25}\mathrm{Mg}$, $^{87}\mathrm{Sr}$, and $^{135,137}\mathrm{Ba}$.
In the latest nuclear quadrupole moment literature Pyykkö (2018), the recommended nuclear quadrupole moment of $^{25}\text{Mg}$ is 0.1994(20) b, as derived from the hyperfine spectrum of the $3s3p\ ^{3}P_{2}$ state Sundholm and Olsen (1991). This value is in good agreement with the 0.201(3) b obtained from mesonic X-ray experiments Weber et al. (1982). Regarding $^{87}\text{Sr}$, the recommended value in Ref. Pyykkö (2018) is 0.305(2) b, which is determined from the hyperfine spectrum of the $\text{Sr}^{+}$ $4d_{5/2}$ state Barwood et al. (2003); Sahoo (2006a). In contrast, Lu et al. reported a value of 0.328(4) b from the $5s5p\ ^{3}P_{1,2}$ states of the neutral Sr atom Lu et al. (2019), approximately 8% larger than the recommended value Pyykkö (2018). For $^{137}\text{Ba}$, the recommended value in Ref. Pyykkö (2018) is 0.236(3) b, extracted from the $\text{Ba}^{+}$ $5d_{3/2}$ and $5d_{5/2}$ states Sahoo et al. (2013). This value is about 4% smaller than those obtained from the $6p_{3/2}$ state Silverans et al. (1986); Wendt et al. (1988). Reported $^{25}\text{Mg}$ nuclear quadrupole moment values are relatively consistent, while there are notable discrepancies in $^{87}\text{Sr}$ and $^{137}\text{Ba}$ values across studies. Thus, a comprehensive and in-depth reinvestigation of $^{87}\text{Sr}$ and $^{137}\text{Ba}$ nuclear quadrupole moments is necessary.
The nuclear quadrupole moment can be obtained from hyperfine spectra computations and measurements of neutral atoms or ions. Actually, the hyperfine spectra of $3s3p~^{3}\!P_{1,2}$ in $^{25}\mathrm{Mg}$ Lurio (1962), $5s5p~^{3}\!P_{1,2}$ and $5s4d~^{1}\!D_{2}$ in $^{87}\mathrm{Sr}$ Grundevik et al. (1983); Zu Putlitz (1963); Heider and Brink (1977), and $6s6p~^{1}\!P_{1}$, $6s5d~^{3}\!D_{1,2,3}$, $6s5d~^{1}\!D_{2}$ in $^{135,137}\mathrm{Ba}$ Gustavsson et al. (1979); Schmelling (1974); Wijngaarden and Li (1995) have been precisely measured, and their electric quadrupole hyperfine-structure constants have been determined. These accurately-determined electric quadrupole hyperfine-structure constants are very suitable for extracting the nuclear quadrupole moments of $^{25}\mathrm{Mg}$, $^{87}\mathrm{Sr}$, and $^{135,137}\mathrm{Ba}$ if the electric-field gradients of these states can be calculated with high precision. The present work aims to systematically calculate electric-field gradients of these low-lying states of neutral Mg, Sr, and Ba atoms. Energies and magnetic-dipole hyperfine-structure constants of these low-lying states are also calculated and compared with available experimental and theoretical data to assess the precision of our results.
The structure of this paper is as follows. In Sec. II, we present the formulations of the hyperfine interaction and the theoretical method. Sec. III is dedicated to presenting the numerical results, along with in-depth discussions. Additionally, in this section, we also compare our results with the available experimental and theoretical data. Finally, Sec. IV provides a summary. Unless explicitly specified otherwise, atomic units are employed throughout the paper.
II Theoretical equations
II.1 Hyperfine interaction
The hyperfine interaction is characterized by the interaction between electrons and the electromagnetic multipole moments of the nucleus. The Hamiltonian can be formulated as follows:
$H_{\rm{HFI}}=\sum_{k}{\textbf{\rm{T}}^{(k)}}\cdot\textbf{\rm{N}}^{(k)},$ (1)
where T and N respectively denote the spherical tensor operators of rank $k$ $(k>0)$ in the electronic and nuclear coordinate spaces. According to the parity and angular selection rules, $k$ must be an even number for electric moments and an odd number for magnetic moments. Specifically, $k=1$, $k=2$, and $k=3$ correspond to the magnetic dipole, electric quadrupole, and magnetic octupole hyperfine interactions, respectively. In this work, we only take into account the magnetic dipole and electric quadrupole hyperfine interactions. The matrix element between two hyperfine states can be computed as:
$\left\langle\gamma^{\prime}IJ^{\prime}F^{\prime}M_{F\prime}^{\prime}\left|H_{\mathrm{HFI}}\right|\gamma IJFM_{F}\right\rangle=\delta_{F^{\prime}F}\delta_{M_{F\prime}^{\prime}M_{F}}\times(-1)^{I+J^{\prime}+F}\sum_{k}\left\{\begin{array}[]{lll}F&J&I\\ k&I&J^{\prime}\end{array}\right\}\left\langle\gamma^{\prime}J^{\prime}\left\|T^{(k)}\right\|\gamma J\right\rangle\left\langle I\left\|N^{(k)}\right\|I\right\rangle,$ (2)
where $I$, $J$, and $F$ represent the nuclear spin, the electron orbital, and the total angular momentum, respectively. The state $|\gamma IJFM_{F}\rangle$ is the hyperfine state constructed by coupling a nuclear eigenstate $|IM_{I}\rangle$ with an atomic eigenstate $|\gamma JM_{J}\rangle$, where $\gamma$ represents the remaining electronic quantum numbers. The first-order correction $\Delta{E_{F}^{(1)}}$ of the hyperfine interaction to the energy can be defined as:
$\Delta{E_{F}^{(1)}}=(-1)^{I+J+F}\sum_{k}\left\{\begin{array}[]{lll}F&J&I\\ k&I&J\end{array}\right\}\left\langle\gamma J\left\|T^{(k)}\right\|\gamma J\right\rangle\left\langle I\left\|N^{(k)}\right\|I\right\rangle.$ (3)
The nuclear matrix elements are expressed in terms of conventional nuclear moments as:
$\left\langle I\left\|N^{(1)}\right\|I\right\rangle=\mu\sqrt{\frac{(I+1)(2I+1)}{I}},\\$ (4)
$\left\langle I\left\|N^{(2)}\right\|I\right\rangle=\frac{Q}{2}\sqrt{\frac{(I+1)(2I+1)(2I+3)}{I(2I-1)}},\\$ (5)
where $\mu$ is the nuclear magnetic dipole moment, and $Q$ is the nuclear electric quadrupole moment. When restricted to $k\leq 2$, $\Delta{E_{F}^{(1)}}$ can be parameterized as:
$\Delta{E_{F}^{(1)}}=$ $\frac{A}{2}K+\frac{B}{2}\frac{3K(K+1)-4I(I+1)J(J+1)}{2I(2I-1)2J(2J-1)},$ (6)
where K=F(F+1)I(I+1)J(J+1)K=F(F+1)-I(I+1)-J(J+1), and $A$ and $B$ are the magnetic dipole and electric quadrupole hyperfine-structure (hfs) constants, which are defined as:
$A=\frac{\mu}{I}\frac{\langle\gamma J\|T^{(1)}\|\gamma J\rangle}{\sqrt{J(J+1)(2J+1)}},$ (7)
and
$B=2Q\bigg[\frac{2J(2J-1)}{(2J+1)(2J+2)(2J+3)}\bigg]^{1/2}\langle\gamma J\|T^{(2)}\|\gamma J\rangle,$ (8)
respectively. Here, $T^{(k)}=\sum_{i}{t^{(k)}(\textbf{r}_{i})}$. The single-particle reduced matrix elements of the operators $t^{(1)}$ and $t^{(2)}$ are given by:
$\begin{cases}\langle{\kappa_{a}}\|t^{(1)}\|\kappa_{b}\rangle=-(\kappa_{a}+\kappa_{b})\langle-\kappa_{a}\|C^{(1)}\|\kappa_{b}\rangle\int_{0}^{\infty}{\frac{f_{a}(r)g_{b}(r)+f_{b}(r)g_{a}(r)}{r^{2}}\times{F^{(1)}(r)}dr}\\ \langle{\kappa_{a}}\|t^{(2)}\|\kappa_{b}\rangle=-\langle\kappa_{a}\|C^{(2)}\|\kappa_{b}\rangle\int_{0}^{\infty}{\frac{f_{a}(r)f_{b}(r)+g_{a}(r)g_{b}(r)}{r^{3}}\times{F^{(2)}(r)}dr}\end{cases},$ (9)
and
$\langle\kappa_{a}\parallel C^{(k)}\parallel\kappa_{b}\rangle$ $=(-1)^{j_{a}+\frac{1}{2}}\sqrt{(2j_{a}+1)(2j_{b}+1)}\times\left(\begin{array}[]{ccc}j_{a}&j_{b}&k\\ -\frac{1}{2}&\frac{1}{2}&0\\ \end{array}\right)\Pi(\ell_{a},L,\ell_{b}),$ (12)
where $\Pi(\ell_{a},k,\ell_{b})=1$ when $\ell_{a}+k+\ell_{b}$ is even; otherwise, $\Pi(\ell_{a},k,\ell_{b})=0$. The relativistic angular-momentum quantum number $\kappa=\ell(\ell+1)-j(j+1)-\frac{1}{4}$. $f$ and $g$ are the large and small radial components of the Dirac wave function, respectively. The nuclear distribution function $F^{(k)}(r)$ is defined as:
$F^{(k)}(r)=\begin{cases}(\frac{r}{R_{N}})^{2k+1},&r\leq{R_{N}}\\ 1,&r>R_{N}\end{cases},$ (13)
where $R_{N}=\sqrt{5/3}\langle{r^{2}}\rangle^{1/2}$ is the radius of the sphere, and $\langle{r^{2}}\rangle^{1/2}$ is the charge root-mean-square radius of the nucleus.
Based on the above definitions of the hfs constants, the nuclear electric quadrupole moment $Q$ (in b) can be extracted from the experimental values of the hfs constant $B$ using the formula:
$Q=\frac{B}{234.9648867q},$ (14)
where the hfs constant $B$ is in the unit of MHz, and the electric-field gradient $q$ is defined as
${q}=2\bigg[\frac{J(2J-1)}{(2J+1)(J+1)(2J+3)}\bigg]^{1/2}\langle\gamma J\|T^{(2)}\|\gamma J\rangle.$ (15)
and is expressed in atomic units.
II.2 Configuration interaction plus coupled-cluster method
In the many-electron atomic system, electron-electron correlation encompasses core-core, core-valence, and valence-valence correlations. All three types of correlation effects are of paramount importance for the precise evaluation of hyperfine interaction properties. In the present work, we adopt a hybrid approach that integrates the relativistic configuration interaction with coupled-cluster methods to comprehensively account for these three types of electron-electron correlations. Within the framework of this hybrid methodology, coupled-cluster calculation is employed to depict the core-core and core-valence correlations, while the valence-valence correlation is taken into consideration through a configuration interaction calculation. This hybrid method was originally developed by Safronova et al. Safronova et al. (2009), and subsequently, Dzuba independently conceived a similar approach Dzuba (2014). The method utilized in our work is conceptually similar to theirs. However, we have independently developed the corresponding software package for this hybrid method, therefore there remain a few differences in the handling details compared to their methods Tang (2025).
For alkaline-earth atoms, the equation governing the effective interaction can be formulated as follows:
$\left(\sum_{i=1}^{2}(H_{\rm DF}(r_{i})+\Sigma_{1}(r_{i}))+\frac{1}{r_{12}}+\Sigma_{2}(r_{12})\right)|\gamma{JM}\rangle=E|\gamma{JM}\rangle,$ (16)
where $H_{\rm DF}$ represents the Dirac-Fock Hamiltonian, and $\Sigma_{1}$ and $\Sigma_{2}$ correspond to the one-body and two-body correlation potentials, respectively. In this work, we derive the one-body and two-body correlation potentials via second-order many-body perturbation (MBPT(2)) calculations and coupled-cluster calculations.
The wave function $|\gamma{JM}\rangle$ of the system is represented as a linear combination of configuration wave functions that share the same parity, angular momentum $J$, and magnetic quantum number $M$. Specifically,
$|\gamma{JM}\rangle=\sum_{v\leq w}C_{vw}|{vw}\rangle,$ (17)
where $C_{vw}$ denote the expansion coefficients. The configuration wave function is constructed from single-particle orbitals as follows:
$|{vw}\rangle=\eta_{vw}\sum_{m_{v},m_{w}}\langle j_{v}m_{v},j_{w}m_{w}|JM\rangle a_{v}^{\dagger}a_{w}^{\dagger}|0\rangle.$ (18)
The symmetry factor $\eta_{vw}$ is defined as:
$\eta_{vw}=\begin{cases}\frac{\sqrt{2}}{2},&v=w\\ 1,&v\neq w\end{cases}.$ (19)
By substituting Eq. (17) into Eq. (16) and applying the variational principle, we can obtain a general eigenvalue equation. Solving this eigenvalue equation enables us to determine the energies and wave functions of the system.
Afterwards, one can compute the transition matrix elements of various operators by utilizing the obtained wave functions. During the calculation of these transition matrix elements, we take into account the random phase approximation (RPA), core Brueckner effects, structural radiation, and normalization corrections of all orders Tang et al. (2017); Blundell et al. (1989). For the convenience of description, later we will combine the core Brueckner effects, structural radiation, and normalization corrections and refer to them as the high-order (HO) correction beyond RPA. Moreover, we include the two-particle (TP) interaction up to the second order Safronova et al. (1999); Savukov (2004). A detailed account of the method and the corresponding formulas has been provided in a recent work Tang (2025).
II.3 Computation details
Table 1: The parameters of the Gauss basis set, along with the basis spaces employed in the CC and CI calculations. NN denotes the number of Gauss basis set. $N_{\rm core}$ represents the quantity of atomic core orbitals. $N_{\rm valence}$ and $N_{\rm virtual}$ signify the numbers of valence orbitals and virtual orbitals in the CC calculation, respectively. $N_{\rm CI}$ stands for the number of orbitals utilized to construct the configurations.
$s$ $p$ $d$ $f$ $g$ $h$ $i$ $k$
Mg
$\alpha(\times{10^{2}})$ 0.085 0.085 0.085 0.26 8.6 8.6 8.6
$\beta$ 1.88 1.87 1.89 1.91 2.0 2.0 2.0
N 35 30 30 25 15 15 15
Ncore 1-2 1 0 0 0 0 0
Nvalence 3-6 2-5 1-2 0 0 0 0
Nvirtual 3-28 2-27 1-25 1-22 1-15 1-15 1-15
NCI 3-22 2-21 1-21 1-19 1-15 0 0
Sr
$\alpha(\times{10^{2}})$ 0.055 0.055 0.055 0.15 8.6 8.6 8.6
$\beta$ 1.85 1.84 1.85 1.91 2.0 2.0 2.0
N 40 40 40 30 25 25 25
Ncore 1-4 1-3 1 0 0 0 0
Nvalence 5-8 4-7 2-5 0 0 0
Nvirtual 5-30 4-29 2-27 1-24 1-20 1-20 1-20
NCI 5-23 4-22 2-21 1-20 1-15 0 0
Ba
$\alpha(\times{10^{2}})$ 0.055 0.055 0.055 0.15 8.6 9.6 9.6 9.6
$\beta$ 1.87 1.85 1.86 1.91 2.0 2.2 2.2 2.2
N 45 45 45 35 30 25 25 25
Ncore 1-5 1-4 1-2 0 0 0 0 0
Nvalence 6-9 5-8 3-6 0 0 0 0
Nvirtual 6-30 5-29 3-27 1-24 1-21 1-20 1-20 1-20
NCI 6-24 5-23 3-23 1-20 1-15 1-15 0 0
In the present work, we use a finite basis set consisting of even-tempered Gaussian-type functions to expand the large and small components of the Dirac wave functions as in Refs. Chaudhuri et al. (1999); Tang (2025). The Gaussian-type function has the form:
$G_{i,\kappa}={\aleph}_{i}r^{n_{\kappa}}e^{-\alpha_{i}{r^{2}}},$ (20)
where $\aleph_{i}$ is the normalization factor, $n_{\kappa}=\ell+1$, and $\alpha_{i}=\alpha\beta^{i-1}$.
Table 1 shows the parameters of the Gauss basis set, along with the basis spaces utilized in the coupled-cluster (CC) and configuration interaction (CI) calculations. $N$ represents the number of the Gauss functions. $N_{\rm core}$ indicates the number of core orbitals. $N_{\rm valence}$ and $N_{\rm virtual}$ respectively denote the numbers of valence orbitals and virtual orbitals involved in the CC calculation. $N_{\rm CI}$ stands for the number of orbitals employed to build the configurations. Specifically, single-particle orbitals with an energy lower than 20000 $a.u.$ are designated as virtual orbitals, and those with an energy lower than 500 $a.u.$ are used for constructing configurations. Nevertheless, in the MBPT calculations, the summation is conducted over the entire basis set.
To obtain the uncertainty of the results, we use five distinct methods to construct the one-body and two-body correlation potentials. Method 1 (CI+MBPT) derives these potentials via second-order many-ody perturbation theory. Method 2 (CI+LCCSD) builds them through linear coupled-cluster singles and doubles calculations, and Method 3 (CI+CCSD) determines them via full coupled-cluster singles and doubles calculations, with the latter considering non-linear terms related to single and double excitations of the cluster operator compared to CI+LCCSD. In Method 2 and Method 3, only contributions from single and double excited states are considered. To compensate for the overlooked higher-order correlation effects, a rescaling parameter $\rho_{\kappa}$ is introduced to replace the one-body correlation potential $\Sigma_{1}$ with $\rho_{\kappa}\Sigma_{1}$, and fine-tuning this parameter can bring the calculated energy closer to the experimental energy, a strategy previously used in CI+MBPT calculations  Zhang et al. (2023, 2024). Method 4 (CI+LCCSDs) starts from LCCSD-obtained potentials, keeps two-body potentials constant and applies a rescaling parameter to the one-body potential. Method 5 (CI+CCSDs) does the same but starts from CCSD-obtained potentials. The rescaling parameters $\rho_{\kappa}$ for CI+LCCSDs and CI+CCSDs are listed in Table 2.
Table 2: The values of the rescaling parameters for CI+LCCSDs and CI+CCSDs methods.
Mg Sr Ba
$\rho_{\kappa}$ LCCSDs CCSDs LCCSDs CCSDs LCCSDs CCSDs
$\rho_{-1}$ 1.011 1.035 0.938 0.988 0.924 0.985
$\rho_{1}$ 1.011 1.050 0.972 1.035 0.979 1.024
$\rho_{-2}$ 1.011 1.050 0.972 1.035 0.979 1.024
$\rho_{2}$ 1.00 1.000 0.957 1.045 0.946 1.041
$\rho_{-3}$ 1.00 1.000 0.957 1.045 0.946 1.041
$\rho_{\rm others}$ 1.00 1.000 1.000 1.000 1.000 1.000
III Results and Discussion
III.1 Energy
Table 3 presents the energies (in cm-1) of the low-lying states of Mg, Sr, and Ba atoms. These energies are obtained by using CI+MBPT(2), CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs methods. To enhance the clarity and readability of the tabular data presentation, we have assigned the abbreviations $\rm M_{1}$, $\rm M_{2}$, $\rm M_{3}$, $\rm M_{4}$, and $\rm M_{5}$ to the CI+MBPT(2), CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs methods, respectively. We also compare our calculated results with the experimental values from National Institute of Standards and Technology (NIST) Kramida et al. (2024). The symbol $\Delta_{n}$ represents the absolute difference between the theoretical results obtained by the $\rm M_{n}$ method and the experimental values, i.e., $\Delta_{n}=|E_{\rm M_{n}}-E_{\rm NIST}|$.
From Table 3, for the Mg atom, CI+MBPT, CI+LCCSD, and CI+CCSD all yield relatively accurate results, with differences from measured values less than 200 cm-1. CI+LCCSD has the smallest difference, suggesting a cancellation between unconsidered correlation effects in CI+CCSD and non-linear terms of single and double excitations. As atomic number $Z$ increases for the three atoms, the difference between CI+MBPT results and measured values rises from 200 cm-1 to 2500 cm-1. The difference for CI+LCCSD results also increases but less significantly. CI+CCSD results are relatively stable, with differences within 250 cm-1, possibly due to more complete correlation effects. CI+LCCSDs and CI+CCSDs, the two methods with scaling parameters, have even smaller differences (controlled within 100 cm-1) from measured values. However, they rely on measured values and are useful for subsequent uncertainty evaluation.
Table 3: The energies (in cm-1) of the low-lying states of Mg, Sr, and Ba atoms, calculated in CI+MBPT(2), CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs approximation, are presented. The CI+MBPT(2), CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs methods are denoted as $\rm M_{1}$, $\rm M_{2}$, $\rm M_{3}$, $\rm M_{4}$, and $\rm M_{5}$, respectively. The experimental values are taken from NIST Kramida et al. (2024). $\Delta_{n}$ represents the absolute difference between the theoretical results obtained by the $\rm M_{n}$ method and the experimental values, ie. $\Delta_{n}=E_{\rm M_{n}}-E_{\rm NIST}$
States $E_{\rm M_{1}}$ $E_{\rm M_{2}}$ $E_{\rm M_{3}}$ $E_{\rm M_{4}}$ $E_{\rm M_{5}}$ $E_{\rm NIST}$ Kramida et al. (2024) $\Delta_{1}$ $\Delta_{2}$ $\Delta_{3}$ $\Delta_{4}$ $\Delta_{5}$
Mg
$3s$ $^{1}\!S_{0}$ –182746 –182903 –182801 –182942 –182933 –182939 193 36 137 3 6
3s3p $^{3}\!P^{o}_{0}$ –160948 –161051 –160971 –161081 –161075 –161088 141 37 118 7 13
3s3p $^{3}\!P^{o}_{1}$ –160927 –161032 –160951 –161062 –161056 –161068 141 36 117 6 12
3s3p $^{3}\!P^{o}_{2}$ –160886 –160990 –160910 –161020 –161014 –161027 141 37 118 7 13
3s3p $^{1}\!P^{o}_{1}$ –147712 –147854 –147766 –147881 –147861 –147887 176 34 122 6 26
3s4s $^{3}\!S_{1}$ –141600 –141710 –141635 –141737 –141727 –141741 141 32 106 5 15
3s4s $^{1}\!S_{0}$ –139294 –139405 –139331 –139431 –139422 –139435 141 31 104 4 13
3s4p $^{3}\!P^{o}_{0}$ –134965 –135073 –135001 –135099 –135090 –135098 133 24 96 2 8
3s4p $^{3}\!P^{o}_{1}$ –134962 –135070 –134998 –135096 –135086 –135094 133 25 96 1 8
3s4p $^{3}\!P^{o}_{2}$ –134955 –135063 –134991 –135089 –135080 –135088 133 24 96 2 8
3s4p $^{1}\!P^{o}_{1}$ –133451 –133569 –133496 –133594 –133583 –133592 141 23 96 2 9
Sr
$5s$ $^{1}\!S_{0}$ –136299 –135478 –135019 –134932 –134935 –134897 1401 581 121 35 38
5s4d $^{3}\!D_{1}$ –118249 –117306 –116517 –116788 –116712 –116738 1511 568 222 49 26
5s4d $^{3}\!D_{2}$ –118178 –117239 –116454 –116723 –116647 –116679 1500 561 225 45 32
5s4d $^{3}\!D_{3}$ –118060 –117136 –116355 –116624 –116544 –116578 1482 558 223 45 35
5s4d $^{1}\!D_{2}$ –116063 –115211 –114534 –114748 –114728 –114748 1315 464 213 0 20
5s5p $^{3}\!P^{o}_{0}$ –121464 –120904 –120573 –120549 –120601 –120580 884 324 6 31 21
5s5p $^{3}\!P^{o}_{1}$ –121278 –120722 –120393 –120368 –120419 –120393 885 329 0 25 26
5s5p $^{3}\!P^{o}_{2}$ –120880 –120321 –119995 –119970 –120018 –119999 881 322 4 29 19
5s5p $^{1}\!P^{o}_{1}$ –114345 –113657 –113237 –113300 –113271 –113199 1146 458 39 102 72
Ba
$6s$ $^{1}\!S_{0}$ –125170 –123485 –122889 –122753 –122795 –122721 2449 764 168 32 74
6s5d $^{3}\!D_{1}$ –116127 –114434 –113473 –113694 –113689 –113687 2440 747 214 7 2
6s5d $^{3}\!D_{2}$ –115915 –114242 –113288 –113507 –113500 –113505 2409 737 218 2 6
6s5d $^{3}\!D_{3}$ –115464 –113848 –112906 –113126 –113107 –113124 2340 723 218 1 18
6s5d $^{1}\!D_{2}$ –113574 –112012 –111118 –111324 –111346 –111326 2248 686 207 1 20
6s6p $^{3}\!P^{o}_{0}$ –111981 –110881 –110454 –110428 –110464 –110455 1526 426 1 26 9
6s6p $^{3}\!P^{o}_{1}$ –111613 –110516 –110091 –110065 –110100 –110084 1529 432 6 19 15
6s6p $^{3}\!P^{o}_{2}$ –110714 –109627 –109212 –109180 –109215 –109206 1508 420 6 26 9
6s6p $^{1}\!P^{o}_{1}$ –106551 –105211 –104660 –104754 –104721 –104661 1890 550 1 94 61
III.2 Magnetic dipole hyperfine-structure constant
Table 4 showcases the hyperfine-structure constant $A$ (in MHz) for the $3s3p~^{3}\!P_{1,2}$ states of 25Mg, the $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states of 87Sr, and the $6s5d~^{3}\!D_{1,2,3}$, $6s5d~^{1}\!D_{2}$, and $6s6p~^{1}\!P_{1}$ states of 137Ba. The data of nuclear spin $I$ and nuclear magnetic dipole moment $\mu$ are from Ref. Stone (2005). Similar to Table 3, we also provide the calculation results of five methods. In the CI+MBPT calculation, the transition matrix elements include RPA and TP corrections. The remaining methods include RPA, HO, and TP corrections.
Table 4: Hyperfine-structure constant $A$ (in MHz) of the low-lying states in 25Mg , 87Sr, and 137Ba
States $A_{\rm M_{1}}$ $A_{\rm M_{2}}$ $A_{\rm M_{3}}$ $A_{\rm M_{4}}$ $A_{\rm M5}$ $A_{\rm Final}$
25Mg $(I=5/2,\mu=-0.85545\mu_{N})$
$3s4s$ $^{3}\!S_{1}$ –324 –321 –320 –321 –321 –321(1)
$3s3p$ $^{3}\!P^{o}_{1}$ –146 –143 –143 –143 –143 –143(1)
$3s3p$ $^{3}\!P^{o}_{2}$ –129 –127 –128 –128 –127 –127(1)
87Sr $(I=9/2,\mu=-1.0936030\mu_{N})$
$5s4d$ $^{1}\!D_{2}$ –0.99 –5.28 –5.51 –5.44 –5.36 –5.3(0.3)
$5s5p$ $^{3}\!P^{o}_{1}$ –285 –259 –257 –256 –257 –259(4)
$5s5p$ $^{3}\!P^{o}_{2}$ –235 –213 –211 –210 –210 –213(4)
137Ba $(I=3/2,\mu=0.937365\mu_{N})$
$6s5d$ $^{3}\!D_{1}$ –635 –514 –509 –500 –503 –514(14)
$6s5d$ $^{3}\!D_{2}$ 467 421 417 410 420 421(11)
$6s5d$ $^{1}\!D_{2}$ –134 –85 –81 –81 –84 –85 (4)
$6s5d$ $^{3}\!D_{3}$ 507 457 455 446 452 457(11)
$6s6p$ $^{1}\!P^{o}_{1}$ –151 –112 –110 –111 –111 -112(2)
From Table 4, for the 25Mg atom, the results of five methods are very close, especially the last four. As atomic number $Z$ increases, CI+MBPT results differ significantly from the other four. For instance, in the $5s4d$ $^{1}\!D_{2}$ state of 87Sr, CI+MBPT result is four-fold smaller, and the hfs constant $A$ of this state is two orders lower than those of other states. In the other seven states of 87Sr and 137Ba, discrepancies between CI+MBPT and the other four exceed $10\%$, reaching up to $50\%$ in the $5s4d$ $^{1}\!D_{2}$ state of 137Ba. We find that the results of CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs are highly similar, with differences in $3\%$. As in our recent work  Tang (2025), we use the CI+LCCSD result as the final recommended value. We compare it with the other three methods’ results and take the maximum difference as the uncertainty. The final recommended value and its uncertainty are listed in the last column of Table 4.
Table 5: Comparison between our results and available experimental and theoretical data of the hyperfine-structure constant $A$.
States $A_{\rm Present}$ $A_{\rm Expt.}$ $A_{\rm Ther.}$
25Mg
$3s4s$ $^{3}\!S_{1}$ –321(1) –321.6(1.5) He et al. (2009) –325 He et al. (2009)
$3s3p$ $^{3}\!P^{o}_{1}$ –143(1) –144.945 Lurio (1962) –141.365 Veseth (1987),–146.1 Porsev and Derevianko (2004)
$3s3p$ $^{3}\!P^{o}_{2}$ –127(1) –128.440 Lurio (1962) –126.287 Veseth (1987),–129.7 Porsev and Derevianko (2004)
–127.5 Beloy et al. (2008)
87Sr
$5s4d$ $^{1}\!D_{2}$ –5.3(0.3) –5.5734(4) Grundevik et al. (1983)
$5s5p$ $^{3}\!P^{o}_{1}$ –259(4) –260.083(5) Zu Putlitz (1963) –258.7 Porsev and Derevianko (2004),–256.96 Lu et al. (2019)
$5s5p$ $^{3}\!P^{o}_{2}$ –213(4) –212.765(1) Heider and Brink (1977) –211.4 Porsev and Derevianko (2004),–230.6 Beloy et al. (2008)
–212.86 Lu et al. (2019)
137Ba
$6s5d$ $^{3}\!D_{1}$ –514(14) –520.602(6) Gustavsson et al. (1979) –547 Kozlov and Porsev (1999)
$6s5d$ $^{3}\!D_{2}$ 421(11) 415.841(30) Gustavsson et al. (1979) 405 Kozlov and Porsev (1999)
$6s5d$ $^{1}\!D_{2}$ –85 (4) –82.180(6) Schmelling (1974) –102 Kozlov and Porsev (1999)
$6s5d$ $^{3}\!D_{3}$ 457(11) 456.548(6) Gustavsson et al. (1979) 443 Kozlov and Porsev (1999)
$6s6p$ $^{1}\!P^{o}_{1} -$112(2) -109.50(13) Wijngaarden and Li (1995) –107 Kozlov and Porsev (1999)
Table 5 compares our final recommended values of hyperfine-structure constants $A$ with available measured and theoretical results. We only list the most accurate measured values Lurio (1962); He et al. (2009); Zu Putlitz (1963); Heider and Brink (1977); Grundevik et al. (1983); Gustavsson et al. (1979); Schmelling (1974); Wijngaarden and Li (1995). Considering uncertainty, our final recommended values match the measured ones. Except for the $5s4d$ $^{1}\!D_{2}$ state of 87Sr and the $5s5d$ $^{1}\!D_{2}$ state of 137Ba, the max difference between our recommended and measured results for other states is within 2%, consistent with the 43Ca atom conclusion Tang (2025). This further validates the reliability of the configuration interaction plus coupled- cluster method.
The last column of Table 5 shows other theoretical results. For the $3s4s$ $^{3}\!S_{1}$ state of 25Mg, there’s one available result from CI+all-order calculation He et al. (2009), equivalent to our CI+LCCSD. Their method only considered RPA, while ours includes RPA, HO, and TP corrections. Considering only RPA, our CI+LCCSD result (323.8 MHz) differs from theirs (325 MHz) by about 1 MHz. For $3s3p$ $^{3}\!P^{o}_{1,2}$ states of 25Mg, Veseth’s many-body perturbation theory results Veseth (1987) differ from ours by $1\%-2\%$. The CI+MBPT results by Beloy et al. Beloy et al. (2008) are consistent with ours, with a difference of less than $1$ MHz. The CI+MBPT results by Porsev et al. Porsev and Derevianko (2004) also match our recommended values.
For the 87Sr atom, our final recommended values are very close to the MCDF results of Lu et al. Lu et al. (2019), and the CI+MBPT results of Poserv et al. Porsev and Derevianko (2004), with a difference of less than $1$ MHz. The CI+MBPT results of Beloy et al. Beloy et al. (2008) are very close to our CI+MBPT results. We have not found that the hyperfine parameters of the $5s4d$ $^{1}\!D_{2}$ state have been calculated by other theoretical methods. In contrast, in the 137Ba atom, the agreement between the CI+MBPT results of Kozlov et al.  Kozlov and Porsev (1999) and the experimental results is not as good as that of our results. This indicates that these states of the 137Ba atom are more sensitive to electron correlation effects, and it also shows that the electron correlation effects beyond the second-order many-body perturbation theory are important. In general, the configuration interaction plus coupled-cluster method can accurately calculate the hyperfine-structure properties of alkaline-earth atom systems.
III.3 Nuclear quadrupole moment
Table 6: The electric-field gradients $q$ (in a.u.) of the low-lying states in 25Mg , 87Sr, and 137Ba.
States $q_{\rm M_{1}}$ $q_{\rm M_{2}}$ $q_{\rm M_{3}}$ $q_{\rm M_{4}}$ $q_{\rm M_{5}}$ $q_{\rm Rec.}$
Mg
$3s3p$ $^{3}\!P^{o}_{1}$ –0.1749 –0.1684 –0.1671 –0.1686 –0.1677 –0.1684(13)
$3s3p$ $^{3}\!P^{o}_{2}$ 0.3482 0.3352 0.3326 0.3355 0.3337 0.3352(26)
Sr
$5s4d$ $^{1}\!D_{2}$ 0.7244 0.6971 0.7039 0.6951 0.7106 0.697(14)
$5s5p$ $^{3}\!P^{o}_{1}$ –0.4939 –0.4490 –0.4415 –0.4464 –0.4452 –0.4490(76)
$5s5p$ $^{3}\!P^{o}_{2}$ 0.9359 0.8525 0.8381 0.8476 0.8451 0.853(14)
Ba
$6s5d$ $^{3}\!D_{1}$ 0.3042 0.3100 0.3161 0.3066 0.3194 0.3100(94)
$6s5d$ $^{3}\!D_{2}$ 0.4632 0.4592 0.4678 0.4541 0.4733 0.459(14)
$6s5d$ $^{1}\!D_{2}$ 1.0878 1.0018 1.0241 0.9964 1.0273 1.002(26)
$6s5d$ $^{3}\!D_{3}$ 0.8199 0.8242 0.8401 0.8155 0.8485 0.824(24)
$6s6p$ $^{1}\!P^{o}_{1}$ 1.0976 0.8768 0.8495 0.8718 0.8586 0.877(27)
Table 6 shows electric-field gradients $q$ of low-lying states of Mg, Sr, and Ba atoms. Like Table 5, it lists five methods’ calculation results, final recommended values, and uncertainties. For Mg, CI+MBPT method’s results differ from the other four by about 3.8%, larger than for hfs constants $A$; differences among the other four are within 1%. In Sr’s $5s5p$ $^{3}P^{o}_{1}$ and $5s5p$ $^{3}P^{o}_{2}$ states, CI+MBPT results differ from the other four by about 10%, while the other four differ within 2%, similar to the hfs constant $A$ results in Table 4. For Sr’s $5s4d$ $^{1}\!D_{2}$ state, the electric-field gradient’s electron-correlation trend differs from the hfs constant $A$. CI+MBPT results differ from the other four by about 4%, but for hfs constant $A$, CI+MBPT result is four times smaller. In Ba’s $6s5d$ $^{3}\!D_{1}$, $6s5d$ $^{3}\!D_{2}$, and $6s5d$ $^{3}\!D_{3}$ states, the five methods’ results differ within 3%, and CI+MBPT result differs from the final recommended value within 2%. For Ba’s $6s5d$ $^{1}\!D_{2}$ and $6s6p$ $^{1}\!P^{o}_{1}$ singlet states, CI+MBPT results differ from the final recommended values by 8.5% and 25% respectively, showing importance of correlation effects beyond MBPT(2). The other four methods’ results differ within 3%. Comparing Table 4 and Table 6, there are differences in electron correlation trends between magnetic-dipole and electric-quadrupole hyperfine interactions.
Table 7: The nuclear quadrupole moment $Q$ (in b) of 25Mg , 87Sr, 135Ba and 137Ba.
State $B_{\rm Expt.}$ $Q$
25Mg
$3s3p$ $^{3}\!P^{o}_{1}$ –8.029 Lurio (1962) 0.2029(21)
$3s3p$ $^{3}\!P^{o}_{2}$ 16.009 Lurio (1962) 0.2032(18)
Final result 0.203(2)
87Sr
$5s4d$ $^{1}\!D_{2}$ 55.421(6) Grundevik et al. (1983) 0.3384(62)
$5s5p$ $^{3}\!P^{o}_{1}$ –35.355(6) Zu Putlitz (1963) 0.3351(64)
$5s5p$ $^{3}\!P^{o}_{2}$ 67.215(15) Heider and Brink (1977) 0.3356(67)
Final result 0.336(4)
137Ba
$6s5d$ $^{3}\!D_{1}$ 17.950(8) Gustavsson et al. (1979) 0.2456(74)
$6s5d$ $^{3}\!D_{2}$ 25.756(50) Gustavsson et al. (1979) 0.2400(89)
$6s5d$ $^{1}\!D_{2}$ 59.568(19) Schmelling (1974) 0.2531(65)
$6s5d$ $^{3}\!D_{3}$ 47.302(60) Gustavsson et al. (1979) 0.2447(76)
$6s6p$ $^{1}\!P^{o}_{1}$ 50.09(21) Wijngaarden and Li (1995) 0.2431(83)
Final result 0.246(4)
135Ba
$6s5d$ $^{3}\!D_{1}$ 11.690(8) Gustavsson et al. (1979) 0.1605(48)
$6s5d$ $^{3}\!D_{2}$ 16.522(54) Gustavsson et al. (1979) 0.1531(57)
$6s5d$ $^{1}\!D_{2}$ 38.713(20) Schmelling (1974) 0.1645(42)
$6s5d$ $^{3}\!D_{3}$ 30.800(62) Gustavsson et al. (1979) 0.1590(49)
$6s6p$ $^{1}\!P^{o}_{1}$ 34.01(22) Wijngaarden and Li (1995) 0.1651(56)
Final result 0.161(2)
By combining the electric-field gradients in Table 6 and the measured electric quadrupole hyperfine-structure constants of these low-lying states Lurio (1962); Grundevik et al. (1983); Zu Putlitz (1963); Heider and Brink (1977); Gustavsson et al. (1979); Schmelling (1974); Wijngaarden and Li (1995), we obtain the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba and 137Ba, which are listed in Table 7. In Table 7, we not only present the nuclear quadrupole moments extracted from individual states, but also provide the final recommended values and their uncertainties through statistical averaging of multiple states. The accuracy of these measured hfs constants $B$ is high. Therefore, the uncertainties of the nuclear quadrupole moments we extracted are entirely due to the electric-field gradients.
For 25Mg, the nuclear quadrupole moments extracted from the $3s3p$ $^{3}\!P^{o}_{1}$ and $3s3p$ $^{3}\!P^{o}_{1}$ states are very close to the final recommended value. The differences among the three are less than 0.2%, and the uncertainty of the final recommended value is approximately 1%. For 87Sr, the differences among the nuclear quadrupole moments obtained from $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states are within 1%.
Compared with the Sr atom, the differences among the nuclear quadrupole moments of 137Ba and 135Ba extracted from the $6s5d$ $^{3}\!D_{1}$, $6s5d$ $^{3}\!D_{2}$, and $6s5d$ $^{3}\!D_{3}$ states are more obvious. For 137Ba, the nuclear quadrupole moment extracted from the $6s5d$ $^{1}\!D_{2}$ state is significantly larger than those from other states. The difference between the nuclear quadrupole moments extracted from single states and the final recommended values is within 3%. However, for 135Ba, the nuclear quadrupole moment extracted from the $6s5d$ $^{3}\!D_{2}$ state is significantly smaller than those from the other four states, and is smaller than the final recommended value by about 5%. We observe that in the $6s5d$ $^{3}\!D_{1}$, $6s5d$ $^{1}\!D_{2}$, and $6s5d$ $^{3}\!D_{3}$ states, the ratios of the hfs constants $B$ of 137Ba to 135Ba are $1.5355$, $1.5387$, and $1.5358$ respectively, which are consistent with the measured value of 1.538485(95) by the nuclear-magnetic-resonance techniques Lutz and Oehler (1978). However, the ratio for the $6s5d$ $^{3}\!D_{2}$ state is 1.5589, which is 1.3% larger than the measured value Lutz and Oehler (1978). In the $6s6p$ $^{1}\!P^{o}_{1}$ state, the ratio of the hfs constants $B$ of 137Ba to 135Ba is $1.4728$, deviating from the measured value Lutz and Oehler (1978) by approximately 4%. Compared with the other four states, the measurement precision of the $6s6p$ $^{1}\!P^{o}_{1}$ state is nearly one order of magnitude lower. The ratio of the nuclear quadrupole moments of 137Ba to 135Ba we obtained through our final recommended values is approximately 1.530, which is smaller than the measured value Lutz and Oehler (1978) by about 0.5%.
Table 8: Comparison between our calculated values and available experimental and theoretical data of the nuclear quadrupole moment $Q$ (in b) of 25Mg , 87Sr, and 137Ba.
$Q$ Sources
25Mg
0.203(2) Present, $3s3p$ $^{3}\!P^{o}_{1}$, $3s3p$ $^{3}\!P^{o}_{2}$
0.1994(20) Sundholm and Olsen (1991), $3s3p$ $^{3}\!P^{o}_{2}$
0.201(3) Weber et al. (1982), Muonic X-ray experiment
0.196(4) Schwalm et al. (1977), Nuclear scattering
87Sr
0.336(4) Present, $5s4d$ $^{1}\!D_{2}$, $5s5p$ $^{3}\!P^{o}_{1}$, $5s5p$ $^{3}\!P^{o}_{2}$
0.305(2) Sahoo (2006a), $4d_{5/2}$
0.328(4) Lu et al. (2019), $5s5p$ $^{3}\!P^{o}_{1}$, $5s5p$ $^{3}\!P^{o}_{2}$
0.335(20) Heider and Brink (1977), $5s5p$ $^{3}\!P^{o}_{1}$, $5s5p$ $^{3}\!P^{o}_{2}$
0.327(24) Mårtensson-Pendrill (2002), $5p_{3/2}$
0.323(20) Yu et al. (2004), $5p_{3/2}$
0.3102(19) Chakraborty et al. (2025), $4d_{5/2}$
137Ba
0.246(4) Present, $6s5d$ $^{3}\!D_{1,2,3}$, $6s5d$ $^{1}\!D_{2}$, $6s6p$ $^{1}\!P^{o}_{1}$
0.236(3) Sahoo et al. (2013), $5d_{3/2}$, $5d_{5/2}$
0.246(2) Silverans et al. (1986), $6p_{3/2}$
0.245(4) Wendt et al. (1988), $6p_{3/2}$
0.246(1) Sahoo (2006b), $6p_{3/2}$
0.2344(17) Chakraborty et al. (2025), $5d_{3/2}$$5d_{5/2}$
Table 8 compares our nuclear quadrupole moments with other experimental and theoretical results Sundholm and Olsen (1991); Weber et al. (1982); Schwalm et al. (1977); Sahoo (2006a); Lu et al. (2019); Heider and Brink (1977); Mårtensson-Pendrill (2002); Yu et al. (2004); Chakraborty et al. (2025); Sahoo et al. (2013); Silverans et al. (1986); Wendt et al. (1988). For 25Mg, considering uncertainties, our results agree well with those from the $3s3p$ $^{3}\!P^{o}_{2}$ state via finite-element multi-configuration Dirac-Fock calculation Sundholm and Olsen (1991) and muonic X-ray experiment Weber et al. (1982), while the result from nuclear scattering Schwalm et al. (1977) is slightly smaller. For 87Sr, we list the nuclear quadrupole moments extracted from the $5p_{3/2}$ and $4d_{5/2}$ states of 87Sr+ Mårtensson-Pendrill (2002); Yu et al. (2004); Sahoo (2006a); Chakraborty et al. (2025), and those extracted from the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states of neutral 87Sr Heider and Brink (1977); Lu et al. (2019). Our results are consistent with the nuclear quadrupole moments extracted from the $5p_{3/2}$ state Mårtensson-Pendrill (2002); Yu et al. (2004), as well as those from the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states Heider and Brink (1977); Lu et al. (2019). Notably, the uncertainty of our results is smaller than that of the nuclear quadrupole moment determined from the $5p_{3/2}$ state Mårtensson-Pendrill (2002); Yu et al. (2004), and is comparable to the uncertainty evaluated by Lu et al. using the MCDF method Lu et al. (2019). However, all these results are larger than the nuclear quadrupole moment extracted from the $4d_{5/2}$ state Sahoo et al. (2013), and this difference is similar to what we previously observed in 43Ca+ and 43Ca Tang (2025). Recently, Chakraborty et al. recalculated the electric-field gradient of the $4d_{5/2}$ state using the coupled-cluster singles, doubles, and triples method and redetermined the nuclear quadrupole moment of 87Sr precisely Chakraborty et al. (2025). Their focus on triple excitations leads to high-accuracy electric-field gradients, so the measured hyperfine-structure constants of the 87Sr+ $4d_{5/2}$ state need further confirmation.
For 137Ba, we compare our results with the nuclear quadrupole moments extracted from the $6p_{3/2}$ and $5d_{3/2,5/2}$ states of 137Ba+ Silverans et al. (1986); Wendt et al. (1988); Sahoo (2006b); Sahoo et al. (2013); Chakraborty et al. (2025). In this table, we do not list the nuclear quadrupole moments extracted from the $6p5d$ and $5s5p$ configurations of neutral Ba atoms in earlier works Jackson and Tuan (1964); Gustavsson et al. (1979); Grundevik et al. (1982). This is because the theoretical values of the hyperfine-structure parameters at that time had low precision, resulting in imprecise nuclear quadrupole moments or large uncertainties.
The comparison shows that our result is in excellent agreement with the nuclear quadrupole moments extracted from the $6p_{3/2}$ state of 137Ba+ Silverans et al. (1986); Wendt et al. (1988); Sahoo (2006b). However, there is a deviation of about 4% compared with the results extracted from the $5d_{3/2,5/2}$ states of 137Ba+ Sahoo et al. (2013); Chakraborty et al. (2025). For the $6p_{3/2}$ state, two groups independently carried out the measurement of its electric quadrupole hyperfine-structure constants Wendt et al. (1988); Villemoes et al. (1993). The measurement results they obtained are in very good agreement, and the measurement precision of both groups is better than 1%. Similarly, for the $5d_{3/2,5/2}$ states, two teams independently measured the hyperfine splittings of these two states Silverans et al. (1986); Lewty et al. (2013). The values of the hfs constants $B$ they obtained are highly consistent, and the measurement precision of the $5d_{3/2,5/2}$ states is at least an order of magnitude higher than that of the $6p_{3/2}$ state. If the theoretical framework can accurately provide the electric-field gradient, then by combining the hfs constants $B$ of the $6p_{3/2}$ or $5d_{3/2,5/2}$ states, the nuclear quadrupole moment can be extracted accurately. In the recent work by Chakraborty et al. Chakraborty et al. (2025), they used the coupled-cluster singles, doubles, and triples method to calculate the electric-field gradient of the $5d_{3/2,5/2}$ states of 137Ba+, and obtained an accurate nuclear quadrupole moment of 137Ba. The uncertainty of their result is extremely small. Our result is obtained by statistically averaging the results of five states of the 137Ba atom. Except for the $6s5d$ $^{1}\!D_{2}$ state, the difference between our calculated magnetic-dipole hyperfine-structure constant and the measured value is within 2% as shown in Table 5. Based on this, we believe that our result should be reliable. However, there is still room for improvement in our method. For example, the contribution of the triple-excitation is not currently included in our CC calculation, and the three-body interaction potential is also ignored, which may be one of the reasons for the 4% difference. However, we think that for the nuclear quadrupole moments of 135,137Ba, it still needs to be further verified by other calculations or new mearsurements. Since the situation of 135Ba is the same as that of 137Ba, we only compared 137Ba in Table 8.
IV Conclusion
In this work, we calculated the magnetic dipole hyperfine-structure constants and electric-field gradients for the $3s3p~^{3}\!P_{1,2}$ states of 24Mg, the $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states of 87Sr, and the $6s5d~^{3}\!D_{1,2,3}$, $6s5d~^{1}\!D_{2}$, and $6s6p~^{1}\!P_{1}$ states of 135,137Ba. We utilized five distinct computational methods, namely CI+MBPT, CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs, to derive the final recommended values along with their corresponding uncertainties. Our final recommended values for the magnetic dipole hyperfine-structure constants of these states are fully consistent with the measured results considering uncertainties.
By incorporating the measured electric quadrupole hyperfine-structure constants and our calculated electric-field gradients of these states, we determined the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba, and 137Ba to be 0.203(2) b, 0.336(4) b, 0.161(2) b, and 0.246(4) b, respectively. The nuclear quadrupole moment of 25Mg determined by us is highly consistent with the value obtained from the muonic X-ray experiment. Our calculated nuclear quadrupole moment of 87Sr is consistent not only with the previously extracted values from the $5p_{3/2}$ state of 87Sr+ but also with those from the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states of 87Sr. Moreover, our nuclear quadrupole moments of 135,137Ba are in agreement with the earlier results derived from the $6p_{3/2}$ state of 135,137Ba+. Nevertheless, there are approximately 10% and 4% discrepancies between our nuclear quadrupole moments of 87Sr and 137Ba and the currently adopted values Pyykkö (2018) obtained from the $4d_{5/2}$ state of 87Sr+ and the $5d_{3/2,5/2}$ state of 137Ba+. Since our nuclear quadrupole moment values were obtained through the statistical averaging of multiple states of neutral atoms, we are confident that our final recommended values are both accurate and reliable. Based on a recently high-precision theoretical work Chakraborty et al. (2025), we think that the previous experimental measurement results of the $4d_{5/2}$ state of 87Sr+ need to be further verified. Given that the contribution of triple excitations is not currently taken into account in our coupled-cluster calculations and the three-body interaction potential is also neglected, there is still room for improvement in our method. We also think that it is still meaningful to further verify the nuclear quadrupole moments of 135,137Ba through other theoretical calculations or new measurements.
Acknowledgements.
The work was supported by the National Natural Science Foundation of China under Grant No.12174268.

### v2

@@
-Determination of nuclear quadrupole moments of 25Mg, 87Sr, and 135,137Ba via configuration-interaction plus coupled-cluster approach
+Determination of nuclear quadrupole moments for 25Mg, 87Sr, and 135,137Ba via configuration-interaction combined with a coupled-cluster approach
 Abstract
@@
 $\left\langle\gamma^{\prime}IJ^{\prime}F^{\prime}M_{F\prime}^{\prime}\left|H_{\mathrm{HFI}}\right|\gamma IJFM_{F}\right\rangle=\delta_{F^{\prime}F}\delta_{M_{F\prime}^{\prime}M_{F}}\times(-1)^{I+J^{\prime}+F}\sum_{k}\left\{\begin{array}[]{lll}F&J&I\\ k&I&J^{\prime}\end{array}\right\}\left\langle\gamma^{\prime}J^{\prime}\left\|T^{(k)}\right\|\gamma J\right\rangle\left\langle I\left\|N^{(k)}\right\|I\right\rangle,$ (2)
-where $I$, $J$, and $F$ represent the nuclear spin, the electron orbital, and the total angular momentum, respectively.
+where $I$, $J$, and $F$ represent the nuclear spin, total angular momentum of the electrons, and the total angular momentum, respectively.
 The state $|\gamma IJFM_{F}\rangle$ is the hyperfine state constructed by coupling a nuclear eigenstate $|IM_{I}\rangle$ with an atomic eigenstate $|\gamma JM_{J}\rangle$, where $\gamma$ represents the remaining electronic quantum numbers.
@@
 where $R_{N}=\sqrt{5/3}\langle{r^{2}}\rangle^{1/2}$ is the radius of the sphere, and $\langle{r^{2}}\rangle^{1/2}$ is the charge root-mean-square radius of the nucleus.
+In the present work, the nucleus is modeled as a uniformly magnetized and charged sphere.
 Based on the above definitions of the hfs constants, the nuclear electric quadrupole moment $Q$ (in b) can be extracted from the experimental values of the hfs constant $B$ using the formula:
@@
 Within the framework of this hybrid methodology, coupled-cluster calculation is employed to depict the core-core and core-valence correlations, while the valence-valence correlation is taken into consideration through a configuration interaction calculation.
-This hybrid method was originally developed by Safronova et al.
+The relativistic configuration interaction plus a linearized version of coupled-cluster theory, called the RCI+all-order method, was first developed by Safronova et al.
 Safronova et al.
-(2009), and subsequently, Dzuba independently conceived a similar approach Dzuba (2014).
+(2009); subsequently, Dzuba independently devised a similar approach Dzuba (2014).
+The concept of this hybrid approach can be traced back to the RCI+MBPT combination introduced by Dzuba and co-workers Dzuba et al.
+(1996).
 The method utilized in our work is conceptually similar to theirs.
@@
 In this work, we derive the one-body and two-body correlation potentials via second-order many-body perturbation (MBPT(2)) calculations and coupled-cluster calculations.
+Here we give only the effective equation for the valence electrons; the core electrons are incorporated into the one- and two-body correlation potentials, while the interaction equations for the many-electron system are detailed in Ref.
+Dzuba et al.
+(1996).
 The wave function $|\gamma{JM}\rangle$ of the system is represented as a linear combination of configuration wave functions that share the same parity, angular momentum $J$, and magnetic quantum number $M$.
@@
 (2017); Blundell et al.
+(1989); Dzuba et al.
 (1998).
 For the convenience of description, later we will combine the core Brueckner effects, structural radiation, and normalization corrections and refer to them as the high-order (HO) correction beyond RPA.
-Moreover, we include the two-particle (TP) interaction up to the second order Safronova et al.
+Moreover, we include the two-particle (TP) interaction up to the second order Dzuba et al.
+(1998); Safronova et al.
 (1999); Savukov (2004).
@@
 Nevertheless, in the MBPT calculations, the summation is conducted over the entire basis set.
-To obtain the uncertainty of the results, we use five distinct methods to construct the one-body and two-body correlation potentials.
+The sizes of the Gaussian basis set, virtual orbital space, and configurations quoted in Table 1 are already fully adequate: any further increase in basis set, virtual orbitals or configurations would shift the energies and hyperfine-structure parameters by less than one unit in the last reported digit.
+In short, within the present theoretical framework, the results are absolutely converged in basis set, virtual orbital, and configuration space.
+To obtain the uncertainty of the results, we use five methods to construct the one-body and two-body correlation potentials.
 Method 1 (CI+MBPT) derives these potentials via second-order many-ody perturbation theory.
@@
 The rescaling parameters $\rho_{\kappa}$ for CI+LCCSDs and CI+CCSDs are listed in Table 2.
+Although CI+LCCSDs and CI+CCSDs methods depend on experimental energies, they remain valuable tools.
+They can provide high-lying excitation energies with appreciable accuracy, and by comparing results obtained with and without scaling parameters we can see how sensitive the property is to higher-order correlation, yielding a rough uncertainty estimate.
+Since the present work targets some low-lying states, we mainly use these two methods to assess the uncertainty of hyperfine-interaction parameters.
 Table 2: The values of the rescaling parameters for CI+LCCSDs and CI+CCSDs methods.
@@
 CI+LCCSDs and CI+CCSDs, the two methods with scaling parameters, have even smaller differences (controlled within 100 cm-1) from measured values.
-However, they rely on measured values and are useful for subsequent uncertainty evaluation.
 Table 3: The energies (in cm-1) of the low-lying states of Mg, Sr, and Ba atoms, calculated in CI+MBPT(2), CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs approximation, are presented.
@@
 $3s4s$ $^{3}\!S_{1}$ –321(1) –321.6(1.5) He et al. (2009) –325 He et al. (2009)
-$3s3p$ $^{3}\!P^{o}_{1}$ –143(1) –144.945 Lurio (1962) –141.365 Veseth (1987),–146.1 Porsev and Derevianko (2004)
-$3s3p$ $^{3}\!P^{o}_{2}$ –127(1) –128.440 Lurio (1962) –126.287 Veseth (1987),–129.7 Porsev and Derevianko (2004)
+$3s3p$ $^{3}\!P^{o}_{1}$ –143(1) –144.945(5) Lurio (1962) –141.365 Veseth (1987),–146.1 Porsev and Derevianko (2004)
+$3s3p$ $^{3}\!P^{o}_{2}$ –127(1) –128.440(5) Lurio (1962) –126.287 Veseth (1987),–129.7 Porsev and Derevianko (2004)
 –127.5 Beloy et al. (2008)
@@
 The last column of Table 5 shows other theoretical results.
-For the $3s4s$ $^{3}\!S_{1}$ state of 25Mg, there’s one available result from CI+all-order calculation He et al.
-(2009), equivalent to our CI+LCCSD.
+For the $3s4s$ $^{3}\!S_{1}$ state of 25Mg, there’s one available result from CI+MBPT calculation He et al.
+(2009).
 Their method only considered RPA, while ours includes RPA, HO, and TP corrections.
@@
 In contrast, in the 137Ba atom, the agreement between the CI+MBPT results of Kozlov et al.
-Kozlov and Porsev (1999) and the experimental results is not as good as that of our results.
+Kozlov and Porsev (1999) and the experimental results is not as good as that of our CI+LCCSD results.
 This indicates that these states of the 137Ba atom are more sensitive to electron correlation effects, and it also shows that the electron correlation effects beyond the second-order many-body perturbation theory are important.
@@
 III.3 Nuclear quadrupole moment
-Table 6: The electric-field gradients $q$ (in a.u.) of the low-lying states in 25Mg , 87Sr, and 137Ba.
+Table 6: The electric-field gradients $q$ (in a.u.) of the low-lying states in 25Mg , 87Sr, and 137Ba. Multiply the electric-field gradient values by 234.9648867 to obtain MHz.
 States $q_{\rm M_{1}}$ $q_{\rm M_{2}}$ $q_{\rm M_{3}}$ $q_{\rm M_{4}}$ $q_{\rm M_{5}}$ $q_{\rm Rec.}$
@@
 $6s6p$ $^{1}\!P^{o}_{1}$ 1.0976 0.8768 0.8495 0.8718 0.8586 0.877(27)
-Table 6 shows electric-field gradients $q$ of low-lying states of Mg, Sr, and Ba atoms.
+Table 6 shows electric-field gradients $q$ (in a.u.) of low-lying states of Mg, Sr, and Ba atoms.
+To express the electric-field-gradient values in MHz, multiply them by 234.9648867.
 Like Table 5, it lists five methods’ calculation results, final recommended values, and uncertainties.
+Like the hfs constant $A$, we adopt the CI+LCCSD result as our final recommended value.
+We compare it with the results from the other three methods and take the maximum difference as the uncertainty.
 For Mg, CI+MBPT method’s results differ from the other four by about 3.8%, larger than for hfs constants $A$; differences among the other four are within 1%.
@@
 25Mg
-$3s3p$ $^{3}\!P^{o}_{1}$ –8.029 Lurio (1962) 0.2029(21)
-$3s3p$ $^{3}\!P^{o}_{2}$ 16.009 Lurio (1962) 0.2032(18)
+$3s3p$ $^{3}\!P^{o}_{1}$ –8.029(5) Lurio (1962) 0.2029(21)
+$3s3p$ $^{3}\!P^{o}_{2}$ 16.009(5) Lurio (1962) 0.2032(18)
 Final result 0.203(2)
@@
 (1979); Schmelling (1974); Wijngaarden and Li (1995), we obtain the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba and 137Ba, which are listed in Table 7.
-In Table 7, we not only present the nuclear quadrupole moments extracted from individual states, but also provide the final recommended values and their uncertainties through statistical averaging of multiple states.
+In Table 7, we not only present the nuclear quadrupole moments extracted from individual states, but also provide the final recommended values and their uncertainties.
+The final recommended values and their uncertainties are obtained by the following equations:
+$\begin{cases}Q_{\rm Final}=\frac{\sum_{i}{\frac{Q_{i}}{(\Delta{Q_{i}})^{2}}}}{\sum_{i}{\frac{1}{(\Delta{Q_{i}})^{2}}}}\\ \Delta{Q_{\rm Final}}=\frac{1}{\sqrt{\sum_{i}{\frac{1}{(\Delta{Q_{i}})^{2}}}}}\end{cases},$ (21)
+where $Q_{i}$ and $\Delta{Q_{i}}$ denote the nuclear quadrupole moment and its uncertainty obtained from the $i$-th state, respectively.
 The accuracy of these measured hfs constants $B$ is high.
 Therefore, the uncertainties of the nuclear quadrupole moments we extracted are entirely due to the electric-field gradients.
-For 25Mg, the nuclear quadrupole moments extracted from the $3s3p$ $^{3}\!P^{o}_{1}$ and $3s3p$ $^{3}\!P^{o}_{1}$ states are very close to the final recommended value.
+For 25Mg, the nuclear quadrupole moments extracted from the $3s3p$ $^{3}\!P^{o}_{1}$ and $3s3p$ $^{3}\!P^{o}_{2}$ states are very close to the final recommended value.
 The differences among the three are less than 0.2%, and the uncertainty of the final recommended value is approximately 1%.
@@
 For example, the contribution of the triple-excitation is not currently included in our CC calculation, and the three-body interaction potential is also ignored, which may be one of the reasons for the 4% difference.
-However, we think that for the nuclear quadrupole moments of 135,137Ba, it still needs to be further verified by other calculations or new mearsurements.
+The recent work by Chakraborty et al.
+shows that triple excitations account for $1-2\%$ of the hfs constants $B$ of the $5d_{3/2}$ and $5d_{5/2}$ stetes in Ba+ Chakraborty et al.
+(2026).
+A contribution of this size is expected to be equally relevant for neutral alkaline-earth atoms.
 Since the situation of 135Ba is the same as that of 137Ba, we only compared 137Ba in Table 8.
@@
 In this work, we calculated the magnetic dipole hyperfine-structure constants and electric-field gradients for the $3s3p~^{3}\!P_{1,2}$ states of 24Mg, the $5s4d~^{1}\!D_{2}$ and $5s5p~^{3}\!P_{1,2}$ states of 87Sr, and the $6s5d~^{3}\!D_{1,2,3}$, $6s5d~^{1}\!D_{2}$, and $6s6p~^{1}\!P_{1}$ states of 135,137Ba.
-We utilized five distinct computational methods, namely CI+MBPT, CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs, to derive the final recommended values along with their corresponding uncertainties.
+We utilized five computational methods, namely CI+MBPT, CI+LCCSD, CI+CCSD, CI+LCCSDs, and CI+CCSDs, to derive the final recommended values along with their corresponding uncertainties.
 Our final recommended values for the magnetic dipole hyperfine-structure constants of these states are fully consistent with the measured results considering uncertainties.
-By incorporating the measured electric quadrupole hyperfine-structure constants and our calculated electric-field gradients of these states, we determined the nuclear quadrupole moments of 25Mg, 87Sr, 135Ba, and 137Ba to be 0.203(2) b, 0.336(4) b, 0.161(2) b, and 0.246(4) b, respectively.
-The nuclear quadrupole moment of 25Mg determined by us is highly consistent with the value obtained from the muonic X-ray experiment.
-Our calculated nuclear quadrupole moment of 87Sr is consistent not only with the previously extracted values from the $5p_{3/2}$ state of 87Sr+ but also with those from the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states of 87Sr.
-Moreover, our nuclear quadrupole moments of 135,137Ba are in agreement with the earlier results derived from the $6p_{3/2}$ state of 135,137Ba+.
+By combining the measured electric-quadrupole hyperfine-structure constants with our calculated electric-field gradients for the relevant states, we obtain nuclear quadrupole moments of 0.203(2) b, 0.336(4) b, 0.161(2) b and 0.246(4) b for 25Mg, 87Sr, 135Ba, and 137Ba, respectively.
+Our value for 25Mg agrees with the muonic X-ray result considering the uncertainty.
+The 87Sr quadrupole moment is consistent with earlier extractions from the $5p_{3/2}$ state of 87Sr+ as well as from the the $5s5p$ $^{3}\!P^{o}_{1}$ and $5s5p$ $^{3}\!P^{o}_{2}$ states of 87Sr, and the nuclear quadrupole moments of 135,137Ba agree with those derived from the $6p_{3/2}$ state of 135,137Ba+.
 Nevertheless, there are approximately 10% and 4% discrepancies between our nuclear quadrupole moments of 87Sr and 137Ba and the currently adopted values Pyykkö (2018) obtained from the $4d_{5/2}$ state of 87Sr+ and the $5d_{3/2,5/2}$ state of 137Ba+.
-Since our nuclear quadrupole moment values were obtained through the statistical averaging of multiple states of neutral atoms, we are confident that our final recommended values are both accurate and reliable.
-Based on a recently high-precision theoretical work Chakraborty et al.
-(2025), we think that the previous experimental measurement results of the $4d_{5/2}$ state of 87Sr+ need to be further verified.
-Given that the contribution of triple excitations is not currently taken into account in our coupled-cluster calculations and the three-body interaction potential is also neglected, there is still room for improvement in our method.
-We also think that it is still meaningful to further verify the nuclear quadrupole moments of 135,137Ba through other theoretical calculations or new measurements.
+These discrepancies between neutral-atom- and ion-based extractions remain to be clarified by further experimental and theoretical works.
+The recent work indicates that the triple excitations contribute $1-2\%$ to the electric-quadrupole hyperfine-structure constants of the $d_{3/2}$ and $d_{5/2}$ states in Sr+ and Ba+ Chakraborty et al.
+(2026).
+To enhance further the accuracy of nuclear quadrupole moments derived with the RCI+CC calculations, the triple excitation contributions need to be included.
 Acknowledgements.